# Gemini and Physical World: Large Language Models Can Estimate the Intensity of Earthquake Shaking from Multi-Modal Social Media Posts

### v2

@@ Abstract
This paper presents a novel approach for estimating the ground shaking intensity using real-time social media data and CCTV footage. Employing the Gemini 1.5 Pro’s[reid2024gemini] model, a multi-modal language model, we demonstrate the ability to extract relevant information from unstructured data utilizing generative AI and natural language processing. The model’s output, in the form of Modified Mercalli Intensity (MMI) values, align well with independent observational data. Furthermore, our results suggest that beyond its advanced visual and auditory understanding abilities, Gemini appears to utilize additional sources of knowledge, including a simplified understanding of the general relationship between earthquake magnitude, distance, and MMI intensity, which it presumably acquired during its training, in its reasoning and decision-making processes. These findings raise intriguing questions about the extent of Gemini’s general understanding of the physical world and its phenomena. Gemini’s ability to generate results consistent with established scientific knowledge highlights the potential of LLMs like Gemini in augmenting our understanding of complex physical phenomena such as earthquakes. More specifically, the results of this study highlight the potential of LLMs like Gemini to revolutionize citizen seismology by enabling rapid, effective, and flexible analysis of crowdsourced data from eyewitness accounts for assessing earthquake impact and providing crisis situational awareness. This approach holds a great promise for improving early warning systems, disaster response, and overall resilience in earthquake-prone regions. This study provides a significant step toward harnessing the power of social media and AI for earthquake disaster mitigation.
@@ Results
Figure 2: showcases an example of Gemini’s output for a social media post. This post documents the ground shaking experienced in Boonton, NJ, located approximately 37.6 km from the epicenter of the M4.8 earthquake that occurred on April 5, 2024, in Tewksbury, New Jersey, USA.
@@
Figure 4: Histograms illustrating the distribution of earthquake intensity with respect to epicentral distance for the New Jersey (a) and Oklahoma (b) earthquakes, utilizing USGS DYFI data. Panel (a) presents a 2D histogram, while panel its margins displays 1D histograms. The blue circles represent the estimated mean Modified Mercalli Intensity (MMI) values for individual social media posts, as determined by the Gemini model.
@@
One possible explanation for the model’s higher confidence in CCTV footage could be the richer set of information typically available in videos, compared to the often limited evidence found in social media posts, such as short tweets. However, this could result in a higher variability in the model’s estimates as well, potentially, due to a wider range of observations/evidence available for the analysis and reasoning. We did not observe a strong correlation between the estimated uncertainties and the confidence values provided by Gemini (Supplementary Figure 1). Interestingly, the narrative descriptions of shaking experiences shared by users in social media posts and interviews, although less direct than the captured shaking moments in CCTV footage and reducing the confidence of the model on its estimate, appear to constrain the model’s estimates, resulting in lower overall uncertainties. However, the less deterministic nature of descriptive evidence leads to a wider range of uncertainties compared to the more visually explicit evidence in videos. Factors such as limited views, low-quality audio, and the lack of informative contextual information in some CCTV footage could contribute to the higher uncertainties observed in MMI estimations for this data type. Further investigation is warranted to better understand the observed differences in uncertainty between data types and to explore methods for improving the model’s performance and confidence across all sources of information. This could involve analyzing the specific types of descriptive evidence that contribute to lower uncertainties, as well as developing techniques to extract more contextual information from CCTV footage.
@@ Discussion
Our results suggest that Gemini can estimate ground motion intensity from social media posts and CCTV videos with a comparable variability to the estimates based on conventional felt reports and instrumental measurements. This capability stems from Gemini’s advanced understanding of language, images, video, and audio. While the model effectively retains detailed and relevant information from the input data in most cases, its capabilities, similar to humans, are not without limitations and can be prone to errors due to factors such as low-quality inputs, limited evidence, and the presence of noise. For example, misinterpretations of visual cues or ambiguous language in social media posts can lead to inaccurate estimations. Beyond its information extraction and understanding abilities, Gemini appears to utilize additional sources of knowledge, presumably acquired during its training, in its reasoning and decision-making processes. Visual inspection of the model’s outputs reveals frequent references to epicentral distance, earthquake magnitude, and post comments in the reasoning section, suggesting that Gemini actively incorporates this contextual information, whether provided directly in the input data or through the prompt, during its analysis. To further investigate the influence of provided context on the output results, we conducted two experiments. In the first experiment, we selected two sample inputs: a CCTV video recorded approximately 5 km from the M7.5 Noto earthquake and a tweet from about 68 km away from the Prague earthquake. In the first experiment, multiple queries were performed on each input sample, varying only the value of the epicentral distance in the prompt while keeping the magnitude fixed at its true value. For both cases, Gemini’s estimated MMI value systematically decreased as larger epicentral distances were used in the prompt, aligning with established principles of earthquake physics and empirical relationships in seismology (Figure 7, top). Similar observations were made when varying the magnitude, the second experiment, while keeping the distance fixed (Figure 7, bottom). These results indicate that Gemini might have a knowledge of relationships between earthquakes magnitude, distance to the epicenter, and ground shaking intensity and actively use it in its predictions. However, our observations, from performing a similar test to multiple other examples, suggest that Gemini uses this knowledge in conjunction with the evidence it extracts from the provided data. This is why its upper and lower estimation bounds, i.e. MMI 8.5 and 2 for Video example and MMI 5.5 and 2 for the Tweet example respectively, remains within the range of viable intensity levels based on the observations and evidence within the input data. As an example, even when the prompt states that provided Tweet was posted by a person 1000 km away from a M 0.5 earthquake, a case where all the empirical relationships predict MMI I or not felt intensity, Gemini still outputs the minimum felt intensity level (MMI II) as it can not ignore the provided evidence by the user who clearly mentioned feeling of an earthquake.

### v3

@@ Abstract
This paper presents a novel approach to extract scientifically valuable information about Earth’s physical phenomena from unconventional sources, such as multi-modal social media posts. Employing a state-of-the-art large language model (LLM), Gemini 1.5 Pro[reid2024gemini], we estimate earthquake ground shaking intensity from these unstructured posts. The model’s output, in the form of Modified Mercalli Intensity (MMI) values, aligns well with independent observational data. Furthermore, our results suggest that LLMs, trained on vast internet data, may have developed a unique understanding of physical phenomena. Specifically, Google’s Gemini models demonstrate a simplified understanding of the general relationship between earthquake magnitude, distance, and MMI intensity, accurately describing observational data even though it’s not identical to established models. These findings raise intriguing questions about the extent to which Gemini’s training has led to a broader understanding of the physical world and its phenomena. The ability of Generative AI models like Gemini to generate results consistent with established scientific knowledge highlights their potential to augment our understanding of complex physical phenomena like earthquakes. The flexible and effective approach proposed in this study holds immense potential for enriching our understanding of the impact of physical phenomena and improving resilience during natural disasters. This research is a significant step toward harnessing the power of social media and AI for natural disaster mitigation, opening new avenues for understanding the emerging capabilities of Generative AI and LLMs for scientific applications.
@@ Results
Figure 2: showcases a sample of Gemini’s output for a social media post. This post documents the ground shaking experienced in Boonton, NJ, located approximately 37.6 km from the epicenter of the M4.8 earthquake that occurred on April 5, 2024, in Tewksbury, New Jersey, USA.
@@
Figure 4: Histograms illustrating the distribution of earthquake intensity with respect to epicentral distance for the New Jersey (a) and Oklahoma (b) earthquakes, utilizing USGS DYFI data. Each panel presents a 2D histogram, while its margins display 1D histograms with counts on the axes. The blue circles represent the estimated mean Modified Mercalli Intensity (MMI) values for individual social media posts, as determined by the Gemini model.
@@
One possible explanation for the model’s higher confidence in CCTV footage could be the richer set of information typically available in videos, compared to the often limited evidence found in social media posts, such as short tweets. However, this could result in a higher variability in the model’s estimates as well, potentially, due to a wider range of observations/evidence available for the analysis and reasoning. We did not observe a strong correlation between the estimated uncertainties and the confidence values provided by Gemini (Supplementary Figure 1). Interestingly, the narrative descriptions of shaking experiences shared by users in social media posts and interviews, although less direct than the captured shaking moments in CCTV footage and reducing the confidence of the model on its estimate, appear to constrain the model’s estimates, resulting in lower overall uncertainties comp;ared to estimates from CCTV footages. However, the less deterministic nature of descriptive evidence leads to a wider range of uncertainties compared to the more visually explicit evidence in videos. Factors such as limited views, low-quality audio, and the lack of informative contextual information in some CCTV footage could contribute to the higher uncertainties observed in MMI estimations for this data type. Further investigation is warranted to better understand the observed differences in uncertainty between data types and to explore methods for improving the model’s performance and confidence across all sources of information. This could involve analyzing the specific types of descriptive evidence that contribute to lower uncertainties, as well as developing techniques to extract more contextual information from CCTV footage.
@@ Discussion
Our results suggest that Gemini can estimate ground motion intensity from social media posts and CCTV videos with a comparable variability to the estimates based on conventional felt reports and instrumental measurements. This capability stems from Gemini’s advanced understanding of language, images, video, and audio. While the model effectively retains detailed and relevant information from the input data in most cases, its capabilities, similar to humans, are not without limitations and can be prone to errors due to factors such as low-quality inputs, limited evidence, and the presence of noise. For example, misinterpretations of visual cues or ambiguous language in social media posts can lead to inaccurate estimations. Beyond its information extraction and understanding abilities, Gemini appears to utilize additional sources of knowledge about general magnitude-distance-intensity relationships of earthquakes, presumably acquired during its training, in its reasoning and decision-making processes. Visual inspection of the model’s outputs reveals frequent references to epicentral distance, earthquake magnitude, and post comments in the reasoning section, suggesting that Gemini actively incorporates this contextual information, whether provided directly in the input data or through the prompt, during its analysis. To further investigate the influence of provided context on the output results, we conducted two experiments. In the first experiment, we selected two sample inputs: a CCTV video recorded approximately 5 km from the M7.5 Noto earthquake and a tweet from about 68 km away from the Prague earthquake. In the first experiment, multiple queries were performed on each input sample, varying only the value of the epicentral distance in the prompt while keeping the magnitude fixed at its true value. For both cases, Gemini’s estimated MMI value systematically decreased as larger epicentral distances were used in the prompt, aligning with established principles of earthquake physics and empirical relationships in seismology (Figure 7, top). Similar observations were made when varying the magnitude, the second experiment, while keeping the distance fixed (Figure 7, bottom). These results indicate that Gemini might have a knowledge of relationships between earthquakes magnitude, distance to the epicenter, and ground shaking intensity and actively use it in its predictions. However, our observations, from performing a similar test to multiple other examples, suggest that Gemini uses this knowledge in conjunction with the evidence it extracts from the provided data. This is why its upper and lower estimation bounds, i.e. MMI 8.5 and 2 for Video example and MMI 5.5 and 2 for the Tweet example respectively, remains within the range of viable intensity levels based on the observations and evidence within the input data. As an example, even when the prompt states that provided Tweet was posted by a person 1000 km away from a M 0.5 earthquake, a case where all the empirical relationships predict MMI I or not felt intensity, Gemini still outputs the minimum felt intensity level (MMI II) as it can not ignore the provided evidence by the user who clearly mentioned feeling of an earthquake.